OpenStudy (anonymous):

Find x if y = –2. 8 – 2xy = 24 A. 2003-05-04-00-00_files/i0370000.jpg B. 4 C. 8 D. –4

OpenStudy (anonymous):

okay so ehres what you do

OpenStudy (anonymous):

we have 8 – 2xy = 24 right?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

8 – 2(-2)y = 24

OpenStudy (anonymous):

then we distribute

OpenStudy (anonymous):

8 + 4y = 24

OpenStudy (anonymous):

then subtract 8 from both sides

OpenStudy (anonymous):

4y = 16

OpenStudy (anonymous):

the divide and you'll have y = 4